# h FR-PHENO-2025-02 A general approach to quantum integration of cross sections in high-energy physics

### v1

@@ Abstract
We present universal building blocks for the quantum integration of generic cross sections in high-energy physics. We make use of Fourier Quantum Monte Carlo Integration as implemented in Quantinuum’s Quantum Monte Carlo Integration engine to provide an extendable methodology for generating efficient circuits that can implement generic cross-section calculations, providing a quadratic speed-up in root mean-squared error convergence with respect to classical Monte Carlo integration. We focus on a concrete example of a $1\to 3$ decay process to illustrate our work.
@@ 1 Introduction
Research in high-energy physics (HEP) aims at exploring fundamental physics at tiny scales. To probe these small scales, elementary particles are collided with very high energies in order to produce other particles. The prime example of such an experiment is the Large Hadron Collider (LHC), where protons are currently collided at a centre-of-mass (COM) energy of $\sqrt{s}=13.6\,\text{TeV}$. In this way, properties of known particles can be studied and new particles, such as the Higgs boson [1, 2], can be discovered.
@@
It is therefore particularly interesting to look at whether quantum processing units can provide an alternative to CPU or GPU technology for computing theoretical predictions in HEP. In the last few years, several works have been devoted to exploring the potential of quantum computing for computing theoretical predictions in perturbative theory [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. The quantum analogue of classical MCI is known as Quantum Monte Carlo Integration (QMCI) [18, 19], and this quantum algorithm utilises the Quantum Amplitude Estimation [20] (QAE) algorithm as a key subroutine, providing a quadratic speed up in root mean-squared error (RMSE) convergence with respect to classical MCI.
@@
In Ref. [21], MP investigated the use of QMCI to integrate elementary scattering processes. In this reference, a variant of the original QAE algorithm was used, namely Iterative Quantum Amplitude Estimation [22] (IQAE). This seminal work focused on simplified versions of scattering cross sections and explored both one- and two-dimensional integrals. Following this, Refs. [23, 24] explored the same idea but using Quantum Fourier Iterative Amplitude Estimation [25] for the quantum integration, which relies on the Fourier Quantum Monte Carlo Integration [19] (FQMCI) method (described later in Section 2.3). In Ref. [25], a small-scale (small number of qubits) parametrised quantum circuit (PQC) is trained to prepare a quantum state representing a specific cross-section integrand, and decompose this state into its Fourier series. Each term in the Fourier series is then integrated separately using QAE, in the same manner as in FQMCI. A potential drawback of this method is that it may not scale efficiently to systems with larger numbers of qubits (as would be required for multi-dimensional integrands to be calculated, for example), due to the trainability issues associated with variational methods in quantum machine learning [26, 27, 28, 29, 30, 31]. Indeed, preparing arbitrary probability distributions on a quantum computer is thought to be computationally hard in the general case, and the topic remains an active field of research.
@@
In the present work, we go significantly beyond what has been done to date by presenting a methodology for the computation of generic cross sections in HEP, applicable to arbitrary dimensionality. In particular, the method of quantum integration makes use of the FQMCI method implemented in Quantinuum’s QMCI engine [32] (developed in part by IW). While research using the QMCI engine has so far been focused on financial applications [33], it is particularly interesting to consider other potential fields such as HEP, where the integrals studied possess different challenges to those in finance e.g., high dimensionality, more complex integrands, and non-trivial integral limits.
@@
More concretely, in this article we investigate the computation of one- and two-dimensional cross-section integrations for scattering processes using the QMCI engine. We demonstrate how one can decompose the integrand for an arbitrary, multi-dimensional cross section into products of constituent building blocks, which are single-variable terms. We describe methods for implementing these individual building blocks on a quantum computer, using the in-built functionality of the QMCI engine. We also discuss how kinematical constraints can be applied to the integrals using this in-built functionality.
@@
In Section 2, we begin by defining the problem and explain why numerical MCI is the preferred method for classical cross-section integration. After this, we introduce FQMCI, Quantinuum’s QMCI engine, and discuss how we can use the engine to perform QMCI for generic cross-section calculations. We then discuss in detail some potential state-preparation methods for generating relativistic Breit-Wigner distributions—one of the key building blocks discussed—in Section 3. Some example applications demonstrating the applicability of our methodology are then discussed in Section 4. Finally, Section 5 contains a discussion of the work, an outlook, and concluding remarks.
@@ 2 Definition of the problem and approaches
2.1 Cross sections in high-energy physics
@@
where $E_{a}$ and $E_{b}$ denote the energy of the initial particles, respectively, while $v_{\rm rel}$ is the relative velocity between them. For fixed energies in the initial state, as is the case in this article, $F$ is a real constant, and thus can simply be factored out. The phase-space term $\mathrm{d}\Phi$ encompasses four-momentum conservation ($p_{a}+p_{b}=p_{c}+p_{d}+\cdots$) as well as the real integration over the kinematic variables of the final-state particles ($c,d,\ldots$). Finally, the matrix element $\mathcal{M}$ encodes the transition probabilities between the states $a,b$ and $c,d,\cdots$. The matrix element itself is constructed from complex numbers and spinors, but when it is complex-conjugate squared [as in Eq. 1] it is always a real number. The underlying theory of interactions between the particles is encoded in the matrix element. In particular, when the interaction between the initial and final state is mediated via virtual particles (not necessarily identical to the external ones), so-called propagators will be present in the matrix element. Their general form, arising in $|\mathcal{M}|^{2}$, reads
@@
where $M$ and $\Gamma$ represent the mass and the decay width of the massive particle, respectively. It is worth mentioning that the decay width of an unstable particle is a measure of its decay probability per unit time. It is the inverse of the particle’s lifetime, and in that respect, an important parameter for characterising a particle’s dynamics. For massless particles, the propagator in Eq. 3 reduces to $p(x)=1/x^{2}$, by setting $M=0$. In this case, the variable $x$ is a kinematic invariant which has units of energy/mass squared. This means that the general form of the integral to be performed in a cross-section calculation [as in Eq. 1] reads
@@
where $N_{I}$ and $N_{P}$ are the number of integration variables and propagators, respectively, so that $N_{I}\geq N_{P}$. In the present notation, the variables of integration $x_{i}$ can either be kinematic invariants or angles related to the final-state particles. The number of integration variables scales as $43n-$ 4 for a $2\to n$ scattering process. The number of propagator terms depends on the process considered, which is defined by the initial- and final-state particles. The index set $I$ contains all combinations of all the $N_{I}$ indices, which are denoted by the subset of indices $S_{k}$. In this way, all possible monomials can be represented. The index, $n_{j}$, represents the power of each term, and is a positive integer. The labels $op$ denote all possible internal massive particles i.e., all particles that are neither in the initial nor final state. For example, for the process that will be studied later (see Fig. 8), the W boson is an internal particle. Finally, $\alpha_{j\ldots k}$ are real constants to ensure full generality. There are as many constants as there are monomials. The challenge is thus to integrate a multi-dimensional integral that has a highly peaked structure due to the propagators. [^3] The propagators describe intermediate particles that can become resonant when the invariant mass of the four momentum is close to the mass of the particle. They are therefore sometimes referred to as resonances. Note that the form of Eq. 4 can be made even more general by allowing sums/differences of random variables instead of simply $x_{l}$ in the denominator. However, given that this case only arises at very high multiplicities (at least $1\to 5$ or $2\to 5$), we defer the treatment of this even more general case for future work.
@@
A further complication when computing cross sections is the appearance of event selections in experimental analyses. Typically, the detectors of an experiment cannot physically cover the entire phase space. Therefore, the integration in Eq. 1 has to be restricted to a smaller domain of integration, corresponding to that physically accessible by the experiment. This is done by setting minimal and maximal values (setting the cut values) to observable quantities, such as the rapidity or the transverse momentum of the final particles. Events with particles that do not pass these requirements are not detected. From the theoretical side, these observable quantities depend on the final-state momenta, which themselves depend on the integration variables. The cross section therefore becomes
@@ 2.2 Classical approach
The most common approach to compute integrals such as the ones presented in the previous section is by reverting to numerical MCI techniques. There are several reasons for this: first, as eluded to above, closed forms for the cut functions do not always exist. This means that numerical techniques must be used. Second, while analytical integration is possible for simple scattering processes (e.g., $2\to 2$ or $2\to 3$), with increasing multiplicity the calculations rapidly become intractable. Third, reverting to analytical calculations make things more difficult to automate, as there are typically many processes that are relevant for a given experiment, e.g., $\mathcal{O}(100)$ different processes for the LHC. The LHC in particular also has the drawback that it is a proton-proton collider, meaning that calculations rely on knowledge of the parton distribution functions for the proton, functions that provide the probability for a parton (an elementary particle such as a quark or gluon) to be emitted. These functions are not known from first principles, and must be extracted from data. As they are typically defined in terms of numerical grids, this makes them more naturally embeddable within a numerical framework. Finally, while one is primarily interested in computing the cross section, this quantity is also measured as a function of other experimental observables such as angles between particles or observables related to energies. When using MCI, computing the cross section as a function of other observables comes at no extra cost, as at each evaluation, the weight used for the cross-section evaluation can also be binned in a histogram according to the value of the observable of interest. In this case, while the binning of the histograms is a computationally cheap operation, for analytical calculations this would require additional, more complicated calculations.
@@
In order to fix notation for later use and clarity, we here define the problem of MCI in a more rigorous way. MCI calculates the expectation of a function, $g(x)$, of a continuous random variable, $X$, with PDF, $f_{X}(x)$, defined as
@@
[…]
Indeed, this can extended further to an arbitrary number of dimensions, and with the integration performed over either all or subsets of the given variables. However, for the purposes of this article it will only be necessary to focus on expectations of separable functions of at most both $X$ and $Y$ i.e., products of functions, $g(x,y)=h(x)l(y)$, with expectation
@@ 2.3 Quantum approach
    A quantum state, |$\ket{p}$, encoding a multivariate probability distribution, $f(.)$, in the values of its complex amplitudes is prepared by applying a quantum circuit, $P$, such that
    |$\ket{p}=P\ket{0}_{n}=\sum_{x}\sqrt{f_{X}(x)}\ket{x}.$ (9)
@@ 2.3.1 Fourier Quantum Monte Carlo Integration
It is important to note for what follows that this technique can also be extended to an arbitrary multivariate case i.e., to expectations of products of functions applied of the form $\mathbb{E}[h(X)l(Y)m(Z)...]$. However, for products of more than two this technique is unlikely to be particularly efficient due to the polynomial increase in the numbers of terms required to compute, and the additional increase in size of the circuits involved.
2.3.2 Quantinuum’s Quantum Monte Carlo Integration engine
@@
First, the engine can take as input any circuit, $P$, provided by the user, and then this circuit is straightforwardly passed along the rest of the QMCI pipeline. It also contains an in-built state-preparation library containing circuits that represent some common distributions, such as the uniform and Gaussian distributions. Second, based on the FQMCI method the engine provides protocols to implement efficient Fourier-decomposition circuits, $R$, for a variety of functions applied, particularly corresponding to moments or products of moments of random variables, such as for $g(x)=x^{2},g(x,y)=xy$, etc. Third, the engine contains a number of in-built algorithms for performing QAE, in general optimised for performance. Fourth, the engine is able to implement limits on integral calculations, based on the use of thresholding operations, implemented via the enhanced $P$-builder functionality. Thresholding operations correspond to binary operations based on some threshold value, $V_{th}$, that enable the estimation of $\mathbb{E}\left[X\,\Theta(X\geq V_{th})\right]$, where the indicator function, $\Theta(z)$, with $z$ either a random variable or an expression, is defined as
@@
Fifth, the user is able to perform a QMCI calculation such that the RMSE of the final estimate is upper bounded by a chosen value. Finally, the engine is equipped with a resource mode that can determine the quantum resources (in terms of qubit numbers and gate counts/depths) required to perform a given QMCI calculation to a desired RMSE. The resource calculator outputs either: with a view to running calculations in the NISQ era—where applying two-qubit gates are the main bottleneck for hardware—the total number and depth of $CX$ gates across all circuits, alongside the same information for the largest circuit that is run, and the same information for the total number and depth of all gates; or, with a view to running calculations in the fault-tolerant era with full quantum error correction—where applying $T$ gates will be the main bottleneck for hardware—the total number and depth of $T$ gates across all circuits, alongside the same information for the largest circuit that is run. In both cases, the total number of qubits required for the largest circuit is also output. Overall, these features will be key to later discussions in this article.
We note that how to split the integrand for the expectation value into the product of function applied, $g(x)$, and probability distribution, $f_{X}(x)$, is in general arbitrary. For QMCI applications in finance, which is so far the only application that has been previously studied using the QMCI engine [33], there is a clear notion of calculating some expectation of the random variable for use in a financial calculation, such as the mean, $g(x)=x$, or the second moment, $g(x)=x^{2}$. However, for general applications (i.e., integrands) there is freedom in how to perform this split, and we take advantage of this notion in the next section.

### v2

@@ Abstract
We present universal building blocks for the quantum integration of generic cross sections in high-energy physics. We make use of Fourier quantum Monte Carlo integration (MCI) as implemented in Quantinuum’s quantum MCI engine to provide an extendable methodology for generating efficient circuits that can implement generic cross-section calculations, providing a quadratic speed-up in root mean-squared error convergence with respect to classical MCI. We focus on a concrete example of a $1\to 3$ decay process to illustrate our work.
@@ 1 Introduction
Research in high-energy physics (HEP) aims at exploring fundamental physics at tiny scales. To probe these small scales, elementary particles are collided with very high energies in order to produce other particles. The prime example of such an experiment is the Large Hadron Collider (LHC), where protons are currently collided at a center-of-mass (COM) energy of $\sqrt{s}=13.6\,\text{TeV}$. In this way, properties of known particles can be studied and new particles, such as the Higgs boson [1, 2], can be discovered.
@@
It is therefore particularly interesting to look at whether quantum processing units can provide an alternative to CPU or GPU technology for computing theoretical predictions in HEP. In the last few years, several works have been devoted to exploring the potential of quantum computing for computing theoretical predictions in perturbative theory [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. The quantum analogue of classical MCI is known as Quantum Monte Carlo Integration (QMCI) [18, 19], and this quantum algorithm utilises the Quantum Amplitude Estimation [20] (QAE) algorithm as a key subroutine, providing a quadratic speed-up in root mean-squared error (RMSE) convergence with respect to classical MCI.
@@
In Ref. [21], Mathieu Pellen (M.P.) investigated the use of QMCI to integrate elementary scattering processes. In this reference, a variant of the original QAE algorithm was used, namely Iterative QAE [22] (IQAE). This seminal work focused on simplified versions of scattering cross sections and explored both one- and two-dimensional integrals. Following this, Refs. [23, 24] explored the same idea but using Quantum Fourier Iterative Amplitude Estimation [25] for the quantum integration, which relies on the Fourier QMCI [19] (FQMCI) method (described later in Section˜2.3). In Ref. [25], a small-scale (small number of qubits) parametrised quantum circuit (PQC) is trained to prepare a quantum state representing a specific cross-section integrand, and decompose this state into its Fourier series. Each term in the Fourier series is then integrated separately using QAE, in the same manner as in FQMCI. A potential drawback of this method is that it may not scale efficiently to systems with larger numbers of qubits (as would be required for multi-dimensional integrands to be calculated, for example), due to the trainability issues associated with variational methods in quantum machine learning [26, 27, 28, 29, 30, 31]. Indeed, preparing arbitrary probability distributions on a quantum computer is thought to be computationally hard in the general case, and the topic remains an active field of research.
@@
In the present work, we go significantly beyond what has been done to date by presenting a methodology for the computation of generic cross sections in HEP, applicable to arbitrary dimensionality. In particular, the method of quantum integration makes use of the FQMCI method implemented in Quantinuum’s QMCI engine [32] (developed in part by Ifan Williams [I.W.]). While research using the QMCI engine has so far been focused on financial applications [33], it is particularly interesting to consider other potential fields such as HEP, where the integrals studied possess different challenges to those in finance e.g. high dimensionality, more complex integrands, and non-trivial integral limits.
@@
More concretely, in this article we investigate the computation of one- and two-dimensional cross-section integrations for scattering processes using the QMCI engine when run on a noiseless state-vector simulator. We demonstrate how one can decompose the integrand for an arbitrary, multi-dimensional cross section into products of constituent building blocks, which are single-variable terms. We describe methods for implementing these individual building blocks on a quantum computer, using the in-built functionality of the QMCI engine. We also discuss how kinematical constraints can be applied to the integrals using this in-built functionality. Our findings demonstrate that this important future application of quantum computing is likely to require fully fault-tolerant hardware before it becomes practically useful.
@@
In Section˜2, we begin by defining the problem and explain why numerical MCI is the preferred method for classical cross-section integration. After this, we introduce FQMCI, Quantinuum’s QMCI engine, and discuss how we can use the engine to perform QMCI for generic cross-section calculations. We then discuss in detail some potential state-preparation methods for generating relativistic Breit-Wigner distributions—one of the key building blocks discussed—in Section˜3. Some example applications demonstrating the applicability of our methodology are then discussed in Section˜4. Finally, Section˜5 contains a discussion of the work, an outlook, and concluding remarks.
@@ 2 Definition of the problem and approaches
2.1 Cross sections in HEP
@@
where $E_{a}$ and $E_{b}$ denote the energy of the initial particles, respectively, while $v_{\rm rel}$ is the relative velocity between them. For fixed energies in the initial state, as is the case in this article, $F$ is a real constant, and thus can simply be factored out. The phase-space term $\mathrm{d}\Phi$ encompasses four-momentum conservation ($p_{a}+p_{b}=p_{c}+p_{d}+\cdots$) as well as the real integration over the kinematic variables of the final-state particles ($c,d,\ldots$). Finally, the matrix element $\mathcal{M}$ encodes the transition probabilities between the states $a,b$ and $c,d,\cdots$. The matrix element itself is constructed from complex numbers and spinors, but when it is complex-conjugate squared [as in Eq.˜1] it is always a real number. The underlying theory of interactions between the particles is encoded in the matrix element. In particular, when the interaction between the initial and final state is mediated via virtual particles (not necessarily identical to the external ones), so-called propagators will be present in the matrix element. Their general form, arising in $|\mathcal{M}|^{2}$, reads
@@
where $M$ and $\Gamma$ represent the mass and the decay width of the massive particle, respectively. It is worth mentioning that the decay width of an unstable particle is a measure of its decay probability per unit time. It is the inverse of the particle’s lifetime, and in that respect, an important parameter for characterising a particle’s dynamics. For massless particles, the propagator in Eq.˜3 reduces to $p(x)=1/x^{2}$, by setting $M=0$. In this case, the variable $x$ is a kinematic invariant which has units of energy/mass squared. This means that the general form of the integral to be performed in a cross-section calculation [as in Eq.˜1] reads
@@
where $N_{I}$ and $N_{P}$ are the number of integration variables and propagators, respectively, so that $N_{I}\geq N_{P}$. In the present notation, the variables of integration $x_{i}$ can either be kinematic invariants or angles related to the final-state particles. The number of integration variables scales as 3n43n-4 for a $2\to n$ scattering process. The number of propagator terms depends on the process considered, which is defined by the initial- and final-state particles. The index set $I$ contains all combinations of all the $N_{I}$ indices, which are denoted by the subset of indices $S_{k}$. In this way, all possible monomials can be represented. The index, $n_{j}$, represents the power of each term, and is a positive integer. The labels $op$ denote all possible internal massive particles i.e. all particles that are neither in the initial nor final state. For example, for the process that will be studied later (see Fig.˜8), the W boson is an internal particle. Finally, $\alpha_{j\ldots k}$ are real constants to ensure full generality. There are as many constants as there are monomials. The challenge is thus to integrate a multi-dimensional integral that has a highly peaked structure due to the propagators. [^4] The propagators describe intermediate particles that can become resonant when the invariant mass of the four momentum is close to the mass of the particle. They are therefore sometimes referred to as resonances. Note that the form of Eq.˜4 can be made even more general by allowing sums/differences of random variables instead of simply $x_{l}$ in the denominator. However, given that this case only arises at very high multiplicities (at least $1\to 5$ or $2\to 5$), we defer the treatment of this even more general case for future work.
@@
A further complication when computing cross sections is the appearance of event selections in experimental analyses. Typically, the detectors of an experiment cannot physically cover the entire phase space. Therefore, the integration in Eq.˜1 has to be restricted to a smaller domain of integration, corresponding to that physically accessible by the experiment. This is done by setting minimal and maximal values (setting the cut values) to observable quantities, such as the rapidity or the transverse momentum of the final particles. [^7] The rapidity of a particle is defined as $y=\frac{1}{2}\log\left(\frac{E+p_{\rm L}}{E-p_{\rm L}}\right)$, where $E$ is the energy of the particle and $p_{\rm L}$ its momentum along the beam axis, typically taken as the $z$ axis by convention. With this convention, the transverse momentum is defined as $p_{\rm T}=\sqrt{p_{x}^{2}+p^{2}_{y}}$. Events with particles that do not pass these requirements are not detected. From the theoretical side, these observable quantities depend on the final-state momenta, which themselves depend on the integration variables. The cross section therefore becomes
@@ 2.2 Classical approach
The most common approach to compute integrals such as the ones presented in the previous section is by reverting to numerical MCI techniques. There are several reasons for this: first, as eluded to above, closed forms for the cut functions do not always exist. This means that numerical techniques must be used. Second, while analytical integration is possible for simple scattering processes (e.g. $2\to 2$ or $2\to 3$), with increasing multiplicity the calculations rapidly become intractable. Third, reverting to analytical calculations make things more difficult to automate, as there are typically many processes that are relevant for a given experiment, e.g. $\mathcal{O}(100)$ different processes for the LHC. The LHC in particular also has the drawback that it is a proton-proton collider, meaning that calculations rely on knowledge of the parton distribution functions for the proton, functions that provide the probability for a parton (an elementary particle such as a quark or gluon) to be emitted. These functions are not known from first principles, and must be extracted from data. As they are typically defined in terms of numerical grids, this makes them more naturally embeddable within a numerical framework. Finally, while one is primarily interested in computing the cross section, this quantity is also measured as a function of other experimental observables such as angles between particles or observables related to energies. When using MCI, computing the cross section as a function of other observables comes at no extra cost, as at each evaluation, the weight used for the cross-section evaluation can also be binned in a histogram according to the value of the observable of interest. In this case, while the binning of the histograms is a computationally cheap operation, for analytical calculations this would require additional, more complicated calculations.
@@
In order to fix notation for later use and clarity, we here define the problem of MCI in a more rigorous way. MCI calculates the expectation of a function, $g(x)$, of a continuous random variable, $X$, with probability density function (PDF), $f_{X}(x)$, defined as
@@
$\mathbb{E}\left[g(X,Y)\right]=\iint g(x,y)f_{XY}(x,y)\,dxdy.$ (7)
Indeed, this can extended further to an arbitrary number of dimensions, and with the integration performed over either all or subsets of the given variables. However, for the purposes of this article it will only be necessary to focus on expectations of separable functions of at most both $X$ and $Y$ i.e. products of functions, $g(x,y)=h(x)l(y)$, with expectation
@@ 2.3 Quantum approach
[…]
@@
At this stage, it is also worth briefly discussing the fact that the first step of QMCI necessarily entails some state-preparation protocol, which has an associated computational cost. A natural question to ask, then, is how the state-preparation overhead affects the overall quadratic advantage of QMCI, and under what assumptions this advantage is fully retained. Herbert [46] discusses this in detail, and notes that the computational advantage is retained if the operational cost of preparing the quantum state encoding the probability distribution is on par with that of generating a classical sample—an efficiency achieved by the $Q$-marginal construction, which allows such quantum states to be prepared directly from reversible classical sampling circuits using only a single layer of Hadamard gates.
@@ 2.3.1 Fourier Quantum Monte Carlo Integration
It is important to note for what follows that this technique can also be extended to an arbitrary multivariate case i.e. to expectations of products of functions applied of the form $\mathbb{E}[h(X)l(Y)m(Z)$...]. However, for products of more than two this technique is unlikely to be particularly efficient due to the polynomial increase in the numbers of terms required to compute, and the additional increase in size of the circuits involved.
2.3.2 Quantinuum’s QMCI engine
@@
First, the engine can take as input any circuit, $P$, provided by the user, and then this circuit is straightforwardly passed along the rest of the QMCI pipeline. It also contains an in-built state-preparation library (the distribution loader) containing circuits that represent some common distributions, such as the uniform and Gaussian distributions (see Section 7 of Ref. [32] for a detailed description of the distribution loader). Second, based on the FQMCI method, the engine provides protocols to implement efficient Fourier-decomposition circuits, $R$, for a variety of functions applied, particularly corresponding to moments or products of moments of random variables, such as for $g(x)=x^{2},g(x,y)=xy$ etc. (see Section 8 of Ref. [32] for a detailed description of the types of quantities that can be estimated using the QMCI engine). Third, the engine contains a number of in-built algorithms for performing QAE, in general optimised for performance [^10] This optimisation generally involves setting the number of shots such that the RMSE of the estimator is minimised. (see Section 9 of Ref [32] for a detailed description of the in-built QAE algorithms). Fourth, the engine is able to implement limits on integral calculations, based on the use of thresholding operations, implemented via the enhanced $P$-builder functionality. Thresholding operations correspond to binary operations based on some threshold value, $V_{th}$, that enable the estimation of $\mathbb{E}\left[X\,\Theta(X\geq V_{th})\right]$, where the indicator function, $\Theta(z)$, with $z$ either a random variable or an expression, is defined as [^11] The threshold is enforced by adding the one’s complement of the classical threshold value, ensuring that the most significant result qubit is set when the condition is met. A subsequent step resets the remaining result qubits to the zero state while preserving the most significant qubit. This makes the cleared qubits available as ancillas for future operations.
@@
Fifth, the user is able to perform a QMCI calculation such that the RMSE of the final estimate is upper bounded by a chosen value (see Section 8 of Ref. [32] for a description of this). Finally, the engine is equipped with a resource mode that can determine the quantum resources (in terms of qubit numbers and gate counts/depths) required to perform a given QMCI calculation to a desired RMSE. The resource calculator outputs either: with a view to running calculations in the noisy intermediate-scale quantum (NISQ) era—where applying two-qubit gates are the main bottleneck for hardware—the total number and depth of $CX$ gates across all circuits, alongside the same information for the largest circuit that is run, and the same information for the total number and depth of all gates; or, with a view to running calculations in the fault-tolerant era with full quantum error correction—where applying $T$ gates will be the main bottleneck for hardware—the total number and depth of $T$ gates across all circuits, alongside the same information for the largest circuit that is run. In both cases, the total number of qubits required for the largest circuit is also output (see Section 10 of Ref. [32] for a detailed description of exactly how the resource mode is implemented). Overall, these features will be key to later discussions in this article.
We note that how to split the integrand for the expectation value into the product of function applied, $g(x)$, and probability distribution, $f_{X}(x)$, is in general arbitrary. For QMCI applications in finance, which is so far the only application that has been previously studied using the QMCI engine [33], there is a clear notion of calculating some expectation of the random variable for use in a financial calculation, such as the mean, $g(x)=x$, or the second moment, $g(x)=x^{2}$. However, for general applications (i.e. integrands) there is freedom in how to perform this split, and we take advantage of this notion in the next section.